OpenStudy (anonymous):

i need help with these problems for my homework step by step 3x^2-x+4x 24x^2=-10x+4 (x-5)(x+4)=2(x-5) 9x^-2x+4=10x

OpenStudy (plainntall):

What do you need to do? Solve for x?

OpenStudy (anonymous):

i have to factor them out

OpenStudy (anonymous):

but im not sure of where to start

OpenStudy (plainntall):

How have you been taught to factor them? Have you learned the quadratic equation yet?

OpenStudy (amtran_bus):

A quick way to check your work would be to plug each "part" or "piece" into a graphing calculator and find the zeros/interceptions.

OpenStudy (anonymous):

i havent been taught yet i was absent last class

OpenStudy (plainntall):

For 3x^2-x+4x can you combine any terms?

OpenStudy (anonymous):

i dont have a graphing calculator, we're not allowed to use them in class

OpenStudy (anonymous):

no you dont combine like terms

OpenStudy (plainntall):

It might help to factor, but what are you suppose to do?

OpenStudy (anonymous):

i dont know what to do thats the problem

OpenStudy (plainntall):

How about you combine the x terms? 3x^2-x+4x

OpenStudy (anonymous):

i think the formula is ax^2+bx+c=0

OpenStudy (plainntall):

3x^2-x+4x=3x^2 + (4x-x)=?

OpenStudy (anonymous):

0 ?

OpenStudy (plainntall):

what is 4x-x=?

OpenStudy (anonymous):

4

OpenStudy (plainntall):

No, you can factor out the x for x(4-1), what is 4-1?

OpenStudy (anonymous):

3

OpenStudy (plainntall):

yes, so 3x^2-x+4x= 3x^2+(4x-x)=3x^2+3 now can you factor anything else out?

OpenStudy (anonymous):

im not sure

OpenStudy (plainntall):

How about the 3? 3x^2+3=3(x^2+1)

OpenStudy (anonymous):

im lost honestly

OpenStudy (plainntall):

Okay Do you understand how to factor things out of equations?

OpenStudy (anonymous):

yes

OpenStudy (plainntall):

If you add, subtract, divide or multiply on one side of an equation you have to do the same on the other side. Sound familiar?

OpenStudy (plainntall):

So x^2+2x=-1 would look like what if you added 1 to both sides?

OpenStudy (plainntall):

Let's do (x-5)(x+4)=2(x-5) What if we divide both sides by (x-5)?